deathscarewolf:

A CD tower is shaped like a cylinder and holds CDs with a diameter of 4.75 inches. If the CD tower is 12 inches tall, what is its volume in cubic inches? Use 3.14 for pi. Round your answer to the nearest hundredth. a. 212.54 cubic inches b. 231.56 cubic inches c. 850.16 cubic inches d. 867.34 cubic inches

jhonyy9:

do you know formula for volume of a cylinder ?

deathscarewolf:

not really

jhonyy9:

V = area of base time high

jhonyy9:

area of base = ?

jhonyy9:

what is the base of a cylinder ?

deathscarewolf:

it would be the 3.14 right?

jhonyy9:

this is just the value of pi

jhonyy9:

do you can draw a cylinder ?

deathscarewolf:

i did before and i just confused my self more

jhonyy9:



jhonyy9:

so how you calcule the volume of this cylinder ?

deathscarewolf:

57?

jhonyy9:

what is this and how you ve got it ?

deathscarewolf:

i timesed 4.75 and 12 to get the 57

jhonyy9:

@Laylalyssa any idea to help ?

deathscarewolf:

i might guess it

Laylalyssa:

So the diameter : 4.75 inches right

jhonyy9:

no the area of base you need multiplie by height

deathscarewolf:

ya

jhonyy9:

@Laylalyssa continue pls

Laylalyssa:

ok so, then radius = 4.65/2 inches.,,right

jhonyy9:

this is the right way

jhonyy9:

4,75

deathscarewolf:

ya

Laylalyssa:

Then the volume of the cylinder would be V= (3.14) (4.75/2)^2 (12)= 212.53875~~ 212.54 in^3. right

deathscarewolf:

yes

Laylalyssa:

There fore, The volume of the cylinder would be

deathscarewolf:

a?

jhonyy9:

@Laylalyssa pls always ste by step

Laylalyssa:

step*

Laylalyssa:

and ok

jhonyy9:

no direct answers

Laylalyssa:

That wasnt the whole answer it was jus part of it-

deathscarewolf:

thx for the help guys

Laylalyssa:

Wait so yk the answer?

deathscarewolf:

a

Laylalyssa:

can I at least tell her if shes correct

jhonyy9:

@deathscarewolf next time you will can solve it alone ?

deathscarewolf:

ya thx for the help you two

Laylalyssa:

ur welcome :)

jhonyy9:

she"s correct than your calcules are right